The position of a particle moving along the x axis is given by (21 22 6 0 )m 2 x = + t − . t , where t is in s.

What is the average velocity during the time interval t = 1.0 s to t = 3.0 s?

a) –6.0 m/s b) –4.0 m/s c) –2.0 m/s d) –8.0 m/s e) 8.0 m/s
Answer #1

X = position of the particle = t-t

Xi = initial position at t =1   ,   = 1-1 = 1

Xf = Final position at t = 3 ,   = 3-3 = 1/27 = 0.037

average velocity = (Xf - Xi)/(3 - 1) = (0.037 - 1)/2 = - 0.482